OpenStudy (anonymous):

Please Help!! A family travels 18 miles down river and returns. It takes 8 hours to make the round trip. Their rate in still water is twice the rate of the rivers current. Find the rate of the current.

OpenStudy (abdullah_me):

18/8=2rate+1rate rate = (18/8)/3=0.75mile/hr

OpenStudy (anonymous):

my answers are A. 1.7 B.9/4 C. 3 D. 6

OpenStudy (anonymous):

@Abdullah_me